# MFC 5.0: An exascale many-physics flow solver

### v1

@@ Abstract
Many problems of interest in engineering, medicine, and the fundamental sciences rely on high-fidelity flow simulation, making performant computational fluid dynamics solvers a mainstay of the open-source software community. A previous work (Bryngelson et al.  Comp. Phys. Comm. (2021)) made MFC 3.0 a published, documented, and open-source source solver with numerous physical features, numerical methods, and scalable infrastructure. MFC 5.0 is a marked update to MFC 3.0, including a broad set of well-established and novel physical models and numerical methods and the introduction of GPU and APU (or superchip) acceleration. We exhibit state-of-the-art performance and ideal scaling on the first two exascale supercomputers, OLCF Frontier and LLNL El Capitan. Combined with MFC’s single-GPU/APU performance, MFC achieves exascale computation in practice. With these capabilities, MFC has evolved into a tool for conducting simulations that many engineering challenge problems hinge upon. New physical features include the immersed boundary method, $N$-fluid phase change, Euler–Euler and Euler–Lagrange sub-grid bubble models, fluid-structure interaction, hypo- and hyper-elastic materials, chemically reacting flow, two-material surface tension, and more. Numerical techniques now represent the current state-of-the-art, including general relaxation characteristic boundary conditions, WENO variants, Strang splitting for stiff sub-grid flow features, and low Mach number treatments. Weak scaling to tens of thousands of GPUs on OLCF Summit and Frontier and LLNL El Capitan see efficiencies within 5% of ideal to over 90% of their respective system sizes. Strong scaling results for a 16-times increase in device count see parallel efficiencies over 90% on OLCF Frontier. Other MFC improvements include ensuring code resilience and correctness with a continuous integration suite, the use of metaprogramming to reduce code length and maintain performance portability, and efficient computational representations for chemical reactions and thermodynamics via code generation with Pyrometheus.
@@ 1.1 The history of MFC
The MFC codebase archaeology starts in the mid-2000s. At this time, MFC was unnamed and developed by Eric Johnsen. This code focused on representing multi-component flow with the diffuse interface method. These developments are described in Johnsen [5]. Following this effort, Vedran Coralic performed a major rewrite and adorned MFC with its name. These features include a true multi-dimensional treatment of the numerical method and modeling, described in Coralic and Colonius [6]. Jomela Meng added treatment for cylindrical coordinates systems and corresponding simulations of high-speed shock–droplet interaction [7]. Kazuki Maeda added a model for bubble dynamics at the sub-grid level via an Euler–Lagrange strategy [8]. Kevin Schmidmayer and Spencer Bryngelson added the thermodynamically consistent 5-equation model of Kapila et al. [2] and the 6-equation model of Saurel et al. [4]. These above efforts were conducted in the research group of Tim Colonius at the California Institute of Technology. Spencer Bryngelson later led a restructuring of MFC at the Georgia Institute of Technology, including additional modeling and numerical features. This effort culminated in the open-sourcing of MFC 3.0 in 2020. The above efforts, in large part, were described in Bryngelson et al. [1]. Modeling, method, and software scalability and portability capabilities were added in subsequent years, which are discussed herein.
1.2 New MFC features
New MFC features include physical models, numerical methods, software resistance through continuous integration and deployment, code coverage testing, and computation acceleration via GPU and APU devices. Physical features include six new models for phase change, non-polytropic sub-grid bubble dynamics, hypo- and hyperelastic materials, chemical reactions and combustion, and surface tension with diffuse interfaces. A ghost-cell immersed boundary method that supports complex geometries with STL and OBJ files was also added. The newly implemented numerical models are general characteristic boundary conditions, improved shock capturing with TENO and WENO-Z constructions, Strang splitting for stiff sub-grid particle dynamics, and special treatments for low-Mach number flows. Additional work on software tools and high-performance computing has made MFC portable and performant on various CPUs, GPUs, and APUs from vendors, including Intel, AMD, and NVIDIA. Significant decreases in run time are also enabled via metaprogramming and static code generation. These additions mark a meaningful step forward in the capabilities and appropriate use cases for MFC.
@@ 1.3 Use of MFC
MFC has a history of being used for early access programs on flagship supercomputers. At the time of writing, these programs include JSC JUPITER (JUREAP, project ExaMFlow) and the LLNL El Capitan and OLCF Frontier final and early access systems, including LLNL Tioga (AMD MI300A testbed) and OLCF Spock and Crusher (AMD MI100 and MI250X testbeds). Commodity cluster use of MFC is broad, including the ACCESS-CI [9] (formerly XSEDE [10]) clusters (PSC Bridges [11] and Bridges2 [12], SDSC Comet [13] and Expanse [14], Purdue Anvil [15], NCSA Delta [16] and DeltaAI [17], TACC Stampede1–3 [18, 19, 20], among others) university clusters, cloud computer systems (Intel’s AI cloud [21] and AMD and Cray’s internal systems, among numerous others). MFC is a SPEChpc benchmark candidate, a benchmarking suite used to evaluate the performance of new and existing supercomputers [22].
@@ 1.4 Manuscript structure
In section 2 we introduce MFC 5.0, describing the code and methods it builds upon while connecting it to MFC 3.0. Section 3 discusses established numerical methods that MFC builds upon. New features follow in section 4, including physical models, numerical methods, and software tooling. We show exemplar MFC simulations in section 5. A discuss other open source flow solvers with some similar features in section 6. Section 7 summarizes the manuscript and MFC 5.0.
@@ 2.1 Multi-component treatment
MFC uses reduced versions of the Baer–Nunziato model [23] and diffuse interface methods to model multiphase flow. Diffuse interface methods model the presence of each species using volume fractions and allow for a small region of smearing between species. Other approaches to multiphase flow modeling require interface tracking, which requires coupled flow solvers and complex mesh management [24]. Alternatively, mesh-free methods like smoothed particle hydrodynamics could be used but smoothed particle hydrodynamics requires extensive load balancing and careful treatment of boundary conditions [24]. Diffuse interface methods are used to avoid these difficulties.
@@ 2.1.1 Five-equation models
The term $K\nabla\cdot\bm{u}$ ensures thermodynamic consistency and accounts for the expansion and compression of each species in mixture regions. The $K\nabla\cdot\bm{u}$ is necessary to model phenomena like spherical bubble collapse [25], but it is non-conservative and can lead to numerical instabilities [25]. For some problems, $K\nabla\cdot\bm{u}$ is not necessary, and the model reduces to that of Allaire et al. [3]. Differences between these models are discussed in Schmidmayer et al. [25].
@@ 2.1.2 Six-equation model
The six-equation model of Saurel et al. [4] allows for pressure disequilibrium between phases and can be used to avoid numerical stability issues associated with the five-equation Kapila model, while maintaining thermodynamic consistency. For two fluids, the six-equation model is
@@ 3.2.2 Approximate Riemann solver
The Riemann problem at each finite volume interface is solved using either the Harten–Lax–van Leer (HLL) or the Harten–Lax-van Leer contact (HLLC) approximate Riemann solver [29]. The HLL approximate Riemann solver admits three constant states separated by two discontinuities with wave speeds $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$. The state at the interface between cell $i$ and $i+1$ is
@@
where
$\bm{q}^{\mathrm{(HLL)}}=\frac{S_{\mathrm{R}}\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}-S_{\mathrm{L}}\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}+\bm{F}(\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k})-\bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k})}{S_{\mathrm{R}}-S_{\mathrm{L}}}$
is the constant start state derived from the integral form of the conservation laws and the consistency condition. The wave speeds $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$ are estimated using the left and right state variables $\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}$ and $\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}$. In the subsonic case, $S_{\mathrm{L}}\leq 0\leq S_{\mathrm{R}}$, the HLL flux is given by
@@ 4.1.1 Immersed boundary method
In MFC 5.0, the solid geometries are parsed. Each cell is characterized either as within a fluid or a solid region. For each immersed boundary, the level set field $\bm{\varphi}$ is calculated. This field stores the vector from each cell to the closest point on the immersed boundary. To enforce the appropriate boundary conditions, the ghost points, image points, and interpolation coefficients are calculated prior to the simulation running with the following algorithm:
The algorithm that applies the boundary conditions follows. MFC first parses the STL or OBJ geometry file (or level set). For each finite volume, we determine if it is in a solid, immersed region via its surface normals. The level set field is computed for each immersed boundary. This field stores the vector from each finite volume center to the closest point on the immersed boundary. For each cell center in the solid regions, if it is within three cells of a fluid cell, we characterize it as a ghost point. For each ghost point, we compute the position of its corresponding image point as $\bm{x}_{\mathrm{(ip)}}=\bm{x}_{\mathrm{(gp)}}+2\bm{\varphi}(\bm{x}_{\mathrm{(gp)}})$, where $\bm{x}_{\mathrm{(ip)}}$ is the image point position, $\bm{x}_{\mathrm{(gp)}}$ is the ghost point position, and $\bm{\varphi}$ is the level set field. The fluid properties at the image points are interpolated from surrounding grid cells, so we compute the interpolation coefficients for each image point using second-order accurate inverse distance weighting. Finite volumes in an immersed solid region are excluded from interpolation.
@@ 4.1.3 Phase change: First order transition, liquid–vapor
$-\sum_{i=3}^{N}\alpha_{i}\rho_{i}q_{i}+$ $[\alpha_{1}\rho_{1}(c_{p,2}-c_{p,1})-\alpha_{m}\rho_{m}c_{p,2}-\sum_{i=3}^{N}\alpha_{i}\rho_{i}c_{p,i}]T=0,$
@@ 4.1.4 Euler–Euler sub-grid bubble dynamics
We model Euler–Euler sub-grid models for bubble dynamics using the method of classes and the method of moments. The method of classes is based on the ensemble phase-averaged equations [37], which can be used to represent dispersions of radially oscillating bubbles as well as other particles [38, 39, 40, 41]. The void fraction of the dispersed phase $\alpha$ is assumed to be negligible compared to the liquid phase ($\alpha_{l}$) under dilute assumptions. The bubbles are represented statistically via random variables $R$, $\dot{R}$, and $R_{o}$ corresponding to the instantaneous bubble radius, time derivative, and equilibrium bubble radius. The mixture-averaged pressure field is
@@
where $N_{R_{o}}$ represents the polydispersity of the bubble population. The conditional moments $\mu_{lm}$ are evaluated using a conditional hyperbolic inversion procedure (CHyQMOM). The CHyQMOM algorithm is described in Fox et al. [44], and a detailed discussion of application to bubble dynamics is in Bryngelson et al. [39]. The total moments follow as
@@ 4.1.6 Fluid–elastic structure interaction
where $\bm{l}$ is the velocity gradient, and for an isotropic Kelvin–Voigt material follows $\dot{\bm{T}}^{e}=2G\dot{\bm{D}}^{d}$ [49] are used to derive a time evolution equation for the elastic stresses. Appending this evolution equation to the five-equation model gives the five-equation model with hypoelasticity

### v3

@@ Abstract
Many problems of interest in engineering, medicine, and the fundamental sciences rely on high-fidelity flow simulation, making performant computational fluid dynamics solvers a mainstay of the open-source software community. A previous work (Bryngelson et al.  Comp. Phys. Comm. (2021)) made MFC 3.0 a published, documented, and open-source solver with numerous physical features, numerical methods, and scalable infrastructure. MFC 5.0 is a marked update to MFC 3.0, including a broad set of well-established and novel physical models and numerical methods and the introduction of GPU and APU (or superchip) acceleration. We exhibit state-of-the-art performance and ideal scaling on the first two exascale supercomputers, OLCF Frontier and LLNL El Capitan. Combined with MFC’s single-accelerator performance, MFC achieves exascale computation in practice. New physical features include the immersed boundary method, $N$-fluid phase change, Euler–Euler and Euler–Lagrange sub-grid bubble models, fluid-structure interaction, hypo- and hyper-elastic materials, chemically reacting flow, two-material surface tension, magnetohydrodynamics (MHD), and more. Numerical techniques now represent the current state-of-the-art, including general relaxation characteristic boundary conditions, WENO variants, Strang splitting for stiff sub-grid flow features, and low Mach number treatments. Weak scaling to tens of thousands of GPUs on OLCF Summit and Frontier and LLNL El Capitan see efficiencies within 5% of ideal to over 90% of their respective system sizes. Strong scaling results for a 16-times increase in device count show parallel efficiencies over 90% on OLCF Frontier. MFC’s software stack has further improved, including continuous integration, which ensures code resilience and correctness through over 300 regression tests; metaprogramming, which reduces code length and maintains performance portability; and code generation for computing chemical reactions.
@@ 1.1 The history of MFC
The MFC codebase archaeology starts in the mid-2000s. At this time, MFC was an unnamed solver developed by Eric Johnsen. This code represented multi-component flow with the diffuse interface method, which were described in Johnsen [5] and Johnsen and Colonius [6]. Following this effort, Vedran Coralic performed a major code rewrite and adorned MFC with its name. These features adorned MFC with a multi-dimensional numerical reconstruction treatment and are described in Coralic and Colonius [7]. Jomela Meng added cylindrical coordinates systems and corresponding simulations of high-speed shock–droplet interaction [8]. Kazuki Maeda added an Euler–Lagrange representation of sub-grid bubble dynamics [9] and acoustic wave generation [10]. Kevin Schmidmayer and Spencer Bryngelson added the 5-equation model of Kapila et al. [2] and the 6-equation model of Saurel et al. [4]. These efforts were conducted in the research group of Tim Colonius at the California Institute of Technology. Spencer Bryngelson later led a restructuring of MFC at the Georgia Institute of Technology, which included additional modeling and numerics features. This effort culminated in the open-sourcing of MFC 3.0 in 2020 and is largely described in Bryngelson et al. [1]. Modeling, method, and software scalability and portability capabilities were added in subsequent years, which are discussed in this manuscript.
1.2 New MFC capabilities
MFC now includes more capable physical models and numerical methods, software resistance through continuous integration and deployment, code coverage testing, and computation acceleration via GPU and APU devices. Physical features include six new models for phase change, non-polytropic sub-grid bubble dynamics, treatment for elastic materials, chemical reactions and combustion, and surface tension. An immersed boundary method that supports complex geometries from analytical level sets or STL files. New numerical methods include generalized characteristic boundary conditions, improved shock capturing with TENO and WENO-Z constructions, Strang splitting for stiff sub-grid dynamics, and treatment for low-Mach number flow. Significant software additions have made MFC a portable and performant solver on CPUs, GPUs, and APUs from various vendors, including Intel, AMD, and NVIDIA. Significant decreases in run time are in part enabled via metaprogramming and static code generation.
@@ 1.3 Use of MFC
MFC has a history of being used for early access programs on flagship supercomputers. At the time of writing, these programs include JSC JUPITER (JUREAP) and the LLNL El Capitan and OLCF Frontier final and early access systems, including LLNL Tioga (AMD MI300A testbed) and OLCF Spock and Crusher (AMD MI100 and MI250X testbeds). Commodity cluster use of MFC is broad, including the ACCESS-CI [11] (formerly XSEDE [12]) clusters (PSC Bridges [13] and Bridges2 [14], SDSC Comet [15] and Expanse [16], Purdue Anvil [17], NCSA Delta [18] and DeltaAI [19], TACC Stampede1–3 [20, 21, 22], among others) university clusters, cloud computer systems (Intel’s AI cloud [23]), and AMD and Cray’s internal systems, among numerous others. MFC is a SPEChpc benchmark candidate, which comprises software that evaluates the performance of supercomputers [24].
@@ 1.4 Manuscript structure
In section 2, we introduce MFC 5.0, describing the code and methods it builds upon while connecting it to MFC 3.0. Section 3 discusses established numerical methods that MFC builds upon. New features follow in section 4, including physical models, numerical methods, and software tooling. We show example MFC simulations in section 5. A discussion of other open source flow solvers with some similar features in section 6. Section 7 summarizes the manuscript and MFC 5.0.
@@ 2.1 Multi-component treatment
MFC uses reduced versions of the Baer–Nunziato model [25] and diffuse interface methods to model multiphase flow. Diffuse interface methods model the presence of each species using volume fractions and allow for a small region of smearing between species. Other approaches to multiphase flow modeling require interface tracking, which requires coupled flow solvers and complex mesh management [26]. Alternatively, mesh-free methods like smoothed particle hydrodynamics could be used, but smoothed particle hydrodynamics requires extensive load balancing and careful treatment of boundary conditions [26]. Diffuse interface methods are used to avoid these difficulties.
@@ 2.1.1 Five-equation models
The term $K\nabla\cdot\bm{u}$ ensures thermodynamic consistency and accounts for the expansion and compression of each species in mixture regions. The $K\nabla\cdot\bm{u}$ is necessary to model phenomena like spherical bubble collapse, but it is non-conservative and can lead to numerical instabilities [27]. For some problems, $K\nabla\cdot\bm{u}$ is not necessary, and the model reduces to that of Allaire et al. [3]. The differences between these models are discussed in Schmidmayer et al. [27].
@@ 2.1.2 Six-equation model
The six-equation model of Saurel et al. [4] allows for pressure disequilibrium between phases and can be used to avoid numerical stability issues associated with the five-equation Kapila model while maintaining thermodynamic consistency. For two fluids, the six-equation model is
@@ 3.2.2 Approximate Riemann solver
The Riemann problem at each finite volume interface is solved using either the Harten–Lax–van Leer (HLL) or the Harten–Lax-van Leer contact (HLLC) approximate Riemann solver [31]. The HLL approximate Riemann solver admits three constant states separated by two discontinuities with wave speeds $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$. The state at the interface between cells $i$ and $i+1$ is
@@
where the constant start state is
$\bm{q}^{\mathrm{(HLL)}}=\frac{S_{\mathrm{R}}\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}-S_{\mathrm{L}}\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}+\bm{F}(\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k})-\bm{F}(\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k})}{S_{\mathrm{R}}-S_{\mathrm{L}}},$
which follows from the integral form of the conservation laws and the consistency condition. The wave speeds $S_{\mathrm{L}}$ and $S_{\mathrm{R}}$ are estimated using the left and right state variables $\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}$ and $\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}$. In the subsonic case, $S_{\mathrm{L}}\leq 0\leq S_{\mathrm{R}}$, the HLL flux is given by
@@ 4.1.1 Immersed boundary method
In MFC 5.0, the solid geometries are parsed. Each cell is characterized either as within a fluid or a solid region. For each immersed boundary, the level set field $\bm{\varphi}$ is calculated. This field stores the vector from each cell to the closest point on the immersed boundary. The ghost points, image points, and interpolation coefficients are calculated and used to enforce the appropriate boundary conditions. These are computed prior to simulation execution and follow the below algorithm.
The algorithm that applies the boundary conditions follows. MFC first parses the STL or OBJ geometry file (or level set). We determine if each finite volume is in a solid, immersed region via its surface normals. The level set field is computed for each immersed boundary. This field stores the vector from each finite volume center to the closest point on the immersed boundary. If each cell center in the solid regions is within three cells of a fluid cell, we characterize it as a ghost point. For each ghost point, we compute the position of its corresponding image point as $\bm{x}_{\mathrm{(ip)}}=\bm{x}_{\mathrm{(gp)}}+2\bm{\varphi}(\bm{x}_{\mathrm{(gp)}})$, where $\bm{x}_{\mathrm{(ip)}}$ is the image point position, $\bm{x}_{\mathrm{(gp)}}$ is the ghost point position, and $\bm{\varphi}$ is the level set field. The fluid properties at the image points are interpolated from surrounding grid cells, so we compute the interpolation coefficients for each image point using second-order accurate inverse distance weighting. Finite volumes in an immersed solid region are excluded from interpolation.
@@ 4.1.3 Phase change: First order transition, liquid–vapor
$-\sum_{i=3}^{N}\alpha_{i}\rho_{i}q_{i}+$ $T\big{[}\alpha_{1}\rho_{1}(c_{p,2}-c_{p,1})-\alpha_{m}\rho_{m}c_{p,2}-\sum_{i=3}^{N}\alpha_{i}\rho_{i}c_{p,i}\big{]}=0,$
@@ 4.1.4 Euler–Euler sub-grid bubble dynamics
We use Euler–Euler sub-grid models for bubble dynamics, including the method of classes and the method of moments. The method of classes is based on the ensemble phase-averaged equations [39], which can be used to represent dispersions of radially oscillating bubbles as well as other particles [40, 41, 42, 43]. The void fraction of the dispersed phase $\alpha$ is assumed to be negligible compared to the liquid phase ($\alpha_{l}$) under dilute assumptions. The bubbles are represented statistically via random variables $R$, $\dot{R}$, and $R_{o}$, corresponding to the instantaneous bubble radius, time derivative, and equilibrium bubble radius. The mixture-averaged pressure field is
@@
where $N_{R_{o}}$ represents the polydispersity of the bubble population. The conditional moments $\mu_{lm}$ are evaluated using a conditional hyperbolic inversion procedure (CHyQMOM). The CHyQMOM algorithm is described in Fox et al. [46], and a detailed discussion of its application to bubble dynamics is in Bryngelson et al. [41]. The total moments follow as
@@ 4.1.6 Fluid–elastic structure interaction
where $\bm{l}$ is the velocity gradient, and for an isotropic Kelvin–Voigt material $\dot{\bm{T}}^{e}=2G\bm{D}^{d}$ [51] are used to derive a time evolution equation for the elastic stresses. Here, the deviatoric component of the velocity gradient is $\bm{D}^{d}=\bm{D}-\mathrm{tr}(\bm{D})\bm{I}/3$. Appending this evolution equation to the five-equation model gives the five-equation model with hypoelasticity
@@ 4.1.8 Surface tension at diffuse material interfaces
4.1.9 Magnetohydrodynamics
Magnetohydrodynamics (MHD) governs the behavior of conducting fluids under the influence of magnetic fields. The governing equations for ideal MHD are [65]
$\frac{\partial\rho}{\partial t}+\nabla\cdot\left(\rho\mathbf{u}\right)$ $=0,$
$\frac{\partial\left(\rho\mathbf{u}\right)}{\partial t}+\nabla\cdot\left[\rho\mathbf{u}\otimes\mathbf{u}+\left(p+\frac{1}{2}|\mathbf{B}|^{2}\right)\mathbf{I}-\mathbf{B}\otimes\mathbf{B}\right]$ $=0,$
$\frac{\partial E}{\partial t}+\nabla\cdot\left[\left(E+p+\frac{1}{2}|\mathbf{B}|^{2}\right)\mathbf{u}-\left(\mathbf{u}\cdot\mathbf{B}\right)\mathbf{B}\right]$ $=0,$
$\frac{\partial\mathbf{B}}{\partial t}+\nabla\cdot\left(\mathbf{u}\otimes\mathbf{B}-\mathbf{B}\otimes\mathbf{u}\right)$ $=0.$
Here, $p$ is the thermal pressure and $\mathbf{B}$ is the magnetic field. The total energy density is $E=\rho e+\frac{1}{2}\rho|\mathbf{u}|^{2}+\frac{1}{2}|\mathbf{B}|^{2}$.
The fast magnetosonic speed, which represents the maximum wave propagation speed, is
$c_{f}=\sqrt{\frac{1}{2}\left(c_{s}^{2}+v_{A}^{2}+\sqrt{\left(c_{s}^{2}+v_{A}^{2}\right)^{2}-4c_{s}^{2}v_{A}^{2}\cos^{2}\theta}\right)},$
where $c_{s}$ is the sound speed and $v_{A}=\sqrt{|\mathbf{B}|^{2}/\rho}$ is the Alfvén speed.
4.1.10 Relativistic Magnetohydrodynamics
Relativistic magnetohydrodynamics (RMHD) extends the classical MHD approach to high-energy density plasmas where fluid velocities approach the speed of light. The governing equations can be expressed in conservative form using the conserved variable vector:
$\mathbf{U}=\begin{pmatrix}D\\[2.5pt] \mathbf{m}\\[2.5pt] \tau\\[2.5pt] \mathbf{B}\end{pmatrix}:=\begin{pmatrix}\Gamma\rho\\[2.5pt] \Gamma^{2}\rho h\mathbf{u}+|\mathbf{B}|^{2}\mathbf{u}-(\mathbf{u}\cdot\mathbf{B})\mathbf{B}\\[2.5pt] \Gamma^{2}\rho h-p+\frac{1}{2}\left[|\mathbf{B}|^{2}+|\mathbf{u}|^{2}|\mathbf{B}|^{2}-(\mathbf{B}\cdot\mathbf{u})^{2}\right]-\Gamma\rho\\[2.5pt] \mathbf{B}\end{pmatrix}.$
Here, $\rho$ denotes the rest-mass density, $\Gamma$ is the Lorentz factor, $p$ is the thermal pressure, and $h=1+e+p/\rho$ is the specific enthalpy. The rest-mass contribution is subtracted from the total energy density to reduce numerical errors at low velocities.
The governing equations are [66]:
$\frac{\partial D}{\partial t}+\nabla\cdot\left(D\mathbf{u}\right)$ $=0,$
$\frac{\partial\mathbf{m}}{\partial t}+\nabla\cdot\left[\mathbf{u}\otimes\mathbf{m}-\frac{1}{\Gamma^{2}}\mathbf{B}\otimes\mathbf{B}-(\mathbf{u}\cdot\mathbf{B})\,\mathbf{u}\otimes\mathbf{B}+p\mathbf{I}\right]$ $=0,$
$\frac{\partial\tau}{\partial t}+\nabla\cdot\left(\mathbf{m}-D\mathbf{u}\right)$ $=0,$
$\frac{\partial\mathbf{B}}{\partial t}+\nabla\cdot\left(\mathbf{u}\otimes\mathbf{B}-\mathbf{B}\otimes\mathbf{u}\right)$ $=0.$
Recovery of the primitive variables from these highly nonlinear conservative variables requires an iterative scheme. In this work, we solve a single nonlinear equation using the Newton-Raphson method [66].